A sequence is bounded if and only if it is both bounded above and bounded

below.


1
Expert's answer
2021-02-19T15:16:36-0500

Let "x_n" be a bounded sequence. "\iff |x_n| \leq K , K\in \mathbb{R} ~ \forall n \in \mathbb{N}"

"\iff -K \leq x_n \leq K"

"\iff x_n" is both bounded above and bounded below.